OpenStudy (anonymous):

solve the equation for x by completing the square. x² = kx + 1

OpenStudy (anonymous):

\[\Large x^2-kx-1=0\] \[\LARGE \Delta=b^2-4ac\] \[\LARGE \Delta=(-k)^2-4\cdot (-1)\] \[\LARGE \Delta=k^2+4\] Now... if: \[\LARGE \Delta>0 \quad \longrightarrow \text{two solutions}\] \[\LARGE \Delta=0 \quad \longrightarrow \text{one solution}\] \[\LARGE \Delta<0 \quad \longrightarrow \text{no solution}\] In our case we have ... ??

OpenStudy (anonymous):

we have two solutions